# Multiplication of 0-1 matrices via clustering

### v1

@@ 5 Extensions
The rows or columns in the input 0-1 matrices can be very long. Also, a large number of clusters might be needed in order to obtain satisfactory low upper bound on their radius. Among other things, for these reasons, we have picked Gonzalez’s classical algorithm for the $k$-center clustering problem [13] as a basic tool in our approach to the arithmetic matrix product of two 0-1 matrices with clustered rows or columns. The running time of his algorithm is linear not only in the number of input points but also in their dimension, and in the parameter $k.$ Importantly, it is very simple and provides a solution within $2$ of the optimum. For instance, there exist faster (in terms of $n$ and $k$) $2$-approximation algorithms for the $k$-center clustering with hidden exponential dependence on the dimension in their running time, see [7, 14],
One can easily generalize our main results by replacing Gonzalez’s algorithm with a simple and efficient approximation algorithm for the more general problem of $k$-center clustering with outliers [3]. In the latter problem, a given number $z$ of input points could be discarded as outliers when trying to minimize the maximum cluster radius. Unfortunately, the algorithms for this more general problem tend to be more complicated and the focus seems to be the approximation ratio achievable in polynomial time (e.g., $3$ in [3] and $2$ in [15]) not the time complexity.
There is a plenty of other variants of $k$-clustering and for the majority of them, plenty of various methods of attacking them, in the literature. In fact, in the design of efficient algorithms for the exact arithmetic matrix product of 0-1 matrices with clustered rows or columns, using the $k$-median clustering could seem more natural. The objective in the latter problem is to minimize the sum of distances between the input points and their nearest centers. Unfortunately, no simple deterministic $O(1)$-approximation algorithms for the latter problem that are efficient in case the dimension and $k$ parameters are large seem be to available [2, 4].

### v2

@@ 5 Extensions
The rows or columns in the input 0-1 matrices can be very long. Also, a large number of clusters might be needed in order to obtain a low upper bound on their radius. Among other things, for these reasons, we have picked Gonzalez’s classical algorithm for the $k$-center clustering problem [16] as a basic tool in our approach to the arithmetic matrix product of two 0-1 matrices with clustered rows or columns. The running time of his algorithm is linear not only in the number of input points but also in their dimension, and in the parameter $k.$ Importantly, it is very simple and provides a solution within $2$ of the optimum. For instance, there exist faster (in terms of $n$ and $k$) $2$-approximation algorithms for $k$-center clustering with hidden exponential dependence on the dimension in their running time, see [10, 17],
One could easily generalize our main results by replacing Gonzalez’s algorithm with a simple and efficient approximation algorithm for the more general problem of $k$-center clustering with outliers [6]. In the latter problem, a given number $z$ of input points could be discarded as outliers when trying to minimize the maximum cluster radius. Unfortunately, the algorithms for this more general problem tend to be more complicated and the focus seems to be the approximation ratio achievable in polynomial time (e.g., $3$ in [6] and $2$ in [18]) not the time complexity.
There are many other variants of clustering than $k$-center clustering, and plenty of methods have been developed for them in the literature. In fact, in the design of efficient algorithms for the exact arithmetic matrix product of 0-1 matrices with clustered rows or columns, using the $k$-median clustering could seem more natural. The objective in the latter problem is to minimize the sum of distances between the input points and their nearest centers. Unfortunately, no simple deterministic $O(1)$-approximation algorithms for the latter problem that are efficient in case the dimension and $k$ parameters are large seem be to available [5, 7].